[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Jan 31, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Jan 31 14:57:25 UTC 2017


Here is today's summary of NHC computer operations:
--- mello

----------

TAFB:
The lx workstation still cannot reboot from the Red Hat menus. We traced  
the problem to hung CAVE processes. Neil Garret and NFC are looking into a  
solution.

Update: missing GFE fields.
The nogaps is ingesting into the Pacific. Not sure why the data was missing  
over the weekend. We will have to monitor.
There are no 6Z or 18Z high resolution non-sbn namnhc runs being sent into  
AWIPS2. We will have to check with Carrissa to see if the data is available.
The gefs and ecens ensemble means will require coordination with the init  
team later today to get the data to ingest into GFE.
.
HSU:
Tuna is still down due an upgrade.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170131/74d16d40/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list